A study comparing the average daily caloric consumption among people on a strict vegan diet, a vegetarian diet, a vegetarian diet plus chicken and fish, and a control group (eating whatever), wit 50 people enrolled in each determine the df values in the following 3 questions.

1. df treatment (diet)=?

2. df error (within)=?

3. df total=?
Answer #1

Solution:

There are total 4 groups.

k = 4

So, df treatment = k - 1 = 4 - 1 = 3

Total number of observations = N = 50

Total df = N - 1 = 50 - 1 = 49

Error df = Total df - Treatment df = 49 - 3 = 46

1. df treatment (diet)=3

2. df error (within)=46

3. df total=49